A vertical diesel engine running at 350 rpm develops 600 kW and has 4 impulses per



revolution. If the fluctuation of energy is 25 per cent of the work done during each



impulse, estimate the cross-sectional area of the rim of the flywheel required to keep the



speed within 2 rpm of the mean speed when the mean peripheral speed of the rim is 1350



m/min. Cast iron has a density of 7.2 Mg/m3



.

1
Expert's answer
2021-11-05T03:29:01-0400

Mean angular velocity, "\omega=350rpm=36.6519rad/s=2199.115rad/min"

"C_s=\frac{2*max\ deviation}{mean\ velocity}"

"=\frac{2(2)}{350}=0.011429"

"r=\frac{mean\ peripheral\ velocity}{\omega}"

"=\frac{1350}{2199.115}=0.61388"

"m=\rho 2πr*A"

"=7200*2\u03c0*0.61388*A=27771.2769A"

"I=mr^2"

"=27771.2769A*0.61388=10465.568A"

"E_m=0.5I\omega^2"

"=0.5*10465.568A*36.6519^2=7029522A"

"i=\frac{350*4}{60}=23.333"

"∆E=\frac{P*0.25}{i}"

"=\frac{600000*0.25}{23.333}=6428.6J"

"\u2206E=2E_mC_s"

"6428.6=2*7029522A*0.011429"

"A=\frac{6428.6}{160680.813}"

=0.0400085m2

"\approx 40000mm^2"